OpenStudy (anonymous):

A sequence is defined by the recursive function f(n + 1) = f(n). If f(3) = 9 , what is f(1) ?

OpenStudy (anonymous):

put n=1,2 f(1+1)=f(1),f(2)=f(1) f(2+1)=f(2) f(3)=f(2)=f(1)=?